OpenStudy (gabbyalicorn):

A volleyball team had 6 boys and 5 girls on it. The team needed 2 equipment managers. The coach selected 1 player at random to be an equipment manager and then selected 1 of the remaining players at random to be the other equipment manager. What is the probability that both equipment managers were girls? Express your answer in simplest form.

OpenStudy (gabbyalicorn):

@freeG13 @gaos

OpenStudy (freeg13):

5 out of 11 are girls then you pick 1 and then pick another 4 out of 10 both are girls

OpenStudy (gabbyalicorn):

u take away 1 from both sides?s

OpenStudy (lulubj):

(5/11)*(4/10) = 2/11

OpenStudy (gabbyalicorn):

but wouldnt it be 20/110

OpenStudy (lulubj):

At first you have 5 girls out of 11 boys/girls on the team. The probability of choosing a girl is 5/11, assuming that we had a success, now we have 4 girls out of 10 boys/girls remaining that can be choosed, now the probability of choosing another girl is 4/10, so... (5/11)*(4/10) = 2/11

OpenStudy (gabbyalicorn):

oooh u simplifyed it

OpenStudy (freeg13):

yes

OpenStudy (gabbyalicorn):

oh kk :) i get it ty :D

OpenStudy (freeg13):

you got it same thing with the last question

OpenStudy (gabbyalicorn):

√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√√